Beth is 14 years older than Aria. Last Year Beth was three times as old as Aria. How Old is Aria?
Answer by ptaylor(2048)

Let x=Aria's age now
Then x-1=Aria's age last year
And x+14=Beth's age now
And (x+14)-1=Beth's age last year
Now we are told that:
(x+14)-1=3(x-1) get rid of parens
x+14-1=3x-3 or
x+13=3x-3 subtract 13 and also 3x from both sides
x-3x+13-13=3x-3x-3-13 collect like terms
-2x=-16 divide both sides by -2
x=8----------------------Aria's age now---------ans
CK
x+14=8+14=22 Beths age now
x+14-1=x+13=8+13=21 Beths age last year
x-1=8-1=7------------------Aria's age last year
Last year Beth was three times as old as Aria
21=7*3
21=21
also
8+14=22
22=22
Hope this helps---ptaylor